OpenStudy (anonymous):

calculate the radius of convergence and the interval of convergence (with examination of the endpoints) \[\sum_{n=1}^{\infty}\frac{3^{n+1}}{n}.x^{n}\] a try... \[|\frac{a_{n+1}}{a_{n}}|=\] \[ \Large \frac{\frac{3^{(n+1)+1}}{n+1}}{\frac{3^{n+1}}{n}}=\] \[\frac{3^{n+2}}{n+1}.\frac{n}{3^{n+1}}=\frac{3}{1}=3\] is it correct and what i need to do additional?

OpenStudy (anonymous):

good work. you are done

OpenStudy (anonymous):

well actually you do have one more step make sure that \(|x|<\frac{1}{3}\)

OpenStudy (anonymous):

i mean you are done, since you get 3 radius of convergence is \(\frac{1}{3}\)

OpenStudy (anonymous):

now replace \(x\) by \(\frac{1}{3}\) and see what you get for the sum. and repeat with \(x=-\frac{1}{3}\)

OpenStudy (anonymous):

I heard there is a definition for that to find its like Limit / 1 but i didnt see, can you post for me a link where definition is there ?

OpenStudy (anonymous):

gotta run, last part should be straightforward, since you are just being asked if a series converges

OpenStudy (anonymous):

ok satellite thank you,